OpenStudy (sarahc):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <1, -2>, v = <-4, 8>

OpenStudy (perl):

take their dot product

OpenStudy (perl):

Let U = (u1,u2) V= (v1,v2) Dot product is U.V = u1 * v1 + u2*v2

OpenStudy (perl):

you also can generalize this Given U = (u1,u2,...,un) V= (v1,v2,...,vn) Dot product is defined as U.V = u1 * v1 + u2*v2 + ... un*vn

OpenStudy (perl):

now there are also some geometric facts since U . V = | U |*|V|*cos (theta) , where theta is between vectors U, V

OpenStudy (perl):

if theta is 90 degrees between the vectors, cos theta = 0, so we have U. V = 0 , when theta is 90 degrees

OpenStudy (perl):

if the angle between them is 180 degrees or 0 degrees, then U and V are parallel . and cos(180) = -1 but its easier to see if two vectors are parallel U is parallel to V when U = k*V

OpenStudy (perl):

u = <1, -2>, v = <-4, 8> i notice that 4*u = v so they are parallel

OpenStudy (sarahc):

@perl thank you very much for your explanation :)